OpenStudy (anonymous):

medal! and i'll fan every one who answers. how much interest is earned on $2,700 in one year if it is invested at 3%? A. $8.10 B. $81.00 C. $810.00

OpenStudy (nurali):

F = P(1+i)^n F = future worth of $ P = present worth of $ i=interest n=years F = 2700(1+0.03)^1 F = 2781 So interest = 2781-2700 = $81

OpenStudy (anonymous):

thanks!

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

@self_made16

OpenStudy (anonymous):

b